Average Senior High School annual cost of tuition fee for all private schools last year was 46,300 a random sample of costs this year for 45 private schools indicated that the sample mean was 47,800 and a sample standard deviation was 5600 at the 0.10 level of significance, is there sufficient evidence to conclude that the cost has increased?


1
Expert's answer
2021-05-07T10:34:42-0400

"\mu = 46300 \\

n = 45 \\

\bar{X} = 47800 \\

\sigma = 5600 \\

z = \frac{\bar{X} - \mu}{\sigma / \sqrt{n}} \\

z = \frac{47800 -46300}{5600 / \sqrt{45}} \\

= \frac{1500}{834.82} \\

= 1.79 \\

Z_{λ/2} = Z_{0.10/2} = Z_{0.05} = 1.645 \\

1.79 > 1.645"

There is sufficient evidence to conclude that the cost has increased.